OpenStudy (anonymous):

The basketball team scored 12 points the first game and 40 points the eighth game of the season. How many points will they earn with this pattern on the 13th game of the season? 58 64 60 62

OpenStudy (anonymous):

The basketball team scored 12 points the first game and 40 points the eighth game of the season. How many points will they earn with this pattern on the 13th game of the season? a = 12 a_8 = 40 difference, d = ( 40 - 12 ) / ( 8 - 1 ) = 28 / 7 .....................= 4 a_n = a + ( n - 1 ) d a_13 = 12 + ( 13 - 1 ) ( 4 ) .........= 12 + ( 12 ) ( 4 ) = 12 + 48 .........= 60

OpenStudy (anonymous):

so the answer is 60

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

:) no problem